# Temporal Test-Time Adaptation with State-Space Models

### v2

@@ 3.2 A Probabilistic Model of Shift Dynamics
where $\phi$ denotes the parameters of the transition and emission density as well as the mixing coefficients, $\phi=\{\psi^{\mathrm{trans}},\psi^{\mathrm{ems}},\bm{\pi}_{1:T}\}$. After one optimization step, we collect the $K$ class prototypes into a matrix ${\mathbf{W}}_{t}$. Using the same hidden representations used to fit ${\mathbf{W}}_{t}$, we generate the predictions using the original predictive model’s softmax parameterization:
$y_{t,n}\ \sim\ \text{{Cat}}\left({\textnormal{y}}_{t,n};\ \text{{softmax}}({\mathbf{W}}_{t}{\mathbf{h}}_{t,n})\right)$ (6)
where $y_{t,n}$ denotes a prediction sampled for the representation vector ${\mathbf{h}}_{t,n}$. Note that adaptation can be performed online by optimizing Eqn. 5 incrementally, considering only up to point $t$. To omit computing the complete-data log likelihood for an increasing sequence as time goes on, we employ a sliding window approach.
@@
The simplest parametric form for the transition and emissions models is Gaussian. The transition noise follows a multivariate Gaussian distribution with zero mean and global covariance $\bm{\Sigma}^{\mathrm{trans}}\in{\mathbb{R}}^{D\times D}$. The resulting model can be seen as a mixture of $K$ Kalman filters (KFs). For posterior inference, thanks to the linearity and Gaussian assumptions, the posterior expectation $\mathbb{E}_{p({\mathbf{W}},{\mathbf{C}}|{\mathbf{H}})}[\cdot]$ in Eqn. 5 can be computed analytically using the well known KF predict, update and smoothing equations (Calabrese & Paninski, 2011; Bishop & Nasrabadi, 2006). However, the closed-form computations come at a cost as they involve inverting matrices of dimensionality $D\times D$. Moreover, the parameter size scales as $K\times D^{2}$, risking overfitting and consuming substantial memory. These are limitations of the Gaussian formulation making it costly for high-dimensional feature spaces and impractical in low resource environments requiring instantaneous predictions. In the next section, we discuss a model for spherical features that circumvents these limitations.
@@ 3.3 Von Mises-Fisher Model for Hyperspherical Features
Choosing Gaussian densities for the transition and emission models, as discussed above, assumes the representation space follows an Euclidean geometry. However, prior work has shown that assuming the hidden representations lie on the unit hypersphere results in a better inductive bias for OOD generalization (Mettes et al., 2019; Bai et al., 2024). This is due to the norms of the representations being biased by in-domain information such as class balance, making angular distances a more reliable signal of class membership in the presence of distribution shift (Mettes et al., 2019; Bai et al., 2024). We too employ the hyperspherical assumption by normalizing the hidden representations such that $||{\mathbf{h}}||_{2}=1$ and model them with the von Mises-Fisher (vMF) distribution (Mardia & Jupp, 2009),
${\text{{vMF}}}({\mathbf{h}};\bm{\mu}_{k},\kappa)=C_{D}(\kappa)\exp\left\{\kappa\cdot\bm{\mu}_{k}^{T}{\mathbf{h}}\right\}$ (7)
where $\bm{\mu}_{k}\in{\mathbb{R}}^{D}$ with $||\bm{\mu}_{k}||_{2}=1$ denotes the mean direction of class $k$, $\kappa\in{\mathbb{R}}^{+}$ the concentration parameter, and $C_{D}(\kappa)$ the normalization constant. High values of $\kappa$ imply larger concentration around $\bm{\mu}_{k}$. The vMF distribution is proportional to a Gaussian distribution with isotropic variance and unit norm. While previous work (Mettes et al., 2019; Ming et al., 2023; Bai et al., 2024) has mainly explored training objectives to encourage latent representations to be vMF-distributed, we apply Eqn. 7 to model the evolving representations.
@@
Returning to the SSM given above, we specify both transition and emission models as vMF distributions, resulting in a hyperspherical transition model, $p({\mathbf{W}}_{t}|{\mathbf{W}}_{t-1})=\prod_{k=1}^{K}{\text{{vMF}}}({\mathbf{w}}_{t,k}|{\mathbf{w}}_{t-1,k},\kappa^{\mathrm{trans}})$, and hyperspherical emission model, $p({\mathbf{H}}_{t}|{\mathbf{W}}_{t})=\prod_{n=1}^{N_{t}}\sum_{k=1}^{K}\pi_{t,k}{\text{{vMF}}}({\mathbf{h}}_{t,n}|{\mathbf{w}}_{t,k},\kappa^{\mathrm{ems}})$. The parameter size of the vMF formulation only scales linearly with the feature dimension, i.e. $\mathcal{O}(DK)$ instead of $\mathcal{O}(D^{2}K)$ as for the Gaussian case. Notably, the noise parameters, $\kappa^{\mathrm{trans}},\kappa^{\mathrm{ems}}$ simplify to scalar values which reduces memory substantially. Fig. 3 illustrates this STAD-vMF variant.
@@
Unlike in the linear Gaussian case, the vMF distribution is not closed under marginalization. Consequentially, the posterior distribution required for the expectation in Eqn. 5, $p({\mathbf{W}}_{1:T}{\mathbf{C}}_{1:T}|{\mathbf{H}}_{1:T})$, cannot be obtained in closed form. We employ a variational EM objective, approximating the posterior with mean-field variational inference, following Gopal & Yang (2014):
$q({\mathbf{w}}_{t,k})={\text{{vMF}}}(\ \cdot\ ;\bm{\rho}_{t,k},\gamma_{t,k})\quad q({\mathbf{c}}_{n,t})=\text{{Cat}}(\ \cdot\ ;\bm{\lambda}_{n,t})\quad\forall t,n,k.$ (8)
The variational distribution $q({\mathbf{W}},{\mathbf{C}})$ factorizes over $n,t,k$ and the objective from Eqn. 5 becomes $\operatorname*{arg\,max}_{\phi}\mathbb{E}_{q({\mathbf{W}},{\mathbf{C}})}\big{[}\log p({\mathbf{H}}_{1:T},{\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T})\big{]}$. More details including the full maximisation steps for $\phi=\{\kappa^{\mathrm{trans}},\kappa^{\mathrm{ems}},\{\{\pi_{t,k}\}_{t=1}^{T}\}_{k=1}^{K}\}$ can be found in Sec. B.1. Notably, posterior inference for the vMF model is much more scalable than the Gaussian case. It operates with linear complexity in $D$, rather than cubic, reducing runtime significantly. Algorithm 1 (Sec. B.2) summarizes the method.
@@
$\begin{split}p&({\textnormal{c}}_{t,n,k}=1|{\mathbf{h}}_{t,n},{\mathbf{w}}_{t,1},\ldots,{\mathbf{w}}_{t,K},\kappa^{\mathrm{ems}})\ =\ \frac{{\text{{vMF}}}({\mathbf{h}}_{t,n};{\mathbf{w}}_{t,k},\kappa^{\mathrm{ems}})}{\sum_{j=1}^{K}{\text{{vMF}}}({\mathbf{h}}_{t,n};{\mathbf{w}}_{t,j},\kappa^{\mathrm{ems}})}\\ &=\frac{C_{D}(\kappa^{\mathrm{ems}})\exp\left\{\kappa^{\mathrm{ems}}\cdot{\mathbf{w}}_{t,k}^{T}{\mathbf{h}}_{t,n}\right\}}{\sum_{j=1}^{K}C_{D}(\kappa^{\mathrm{ems}})\exp\left\{\kappa^{\mathrm{ems}}\cdot{\mathbf{w}}_{t,j}^{T}{\mathbf{h}}_{t,n}\right\}}\ =\ \text{{softmax}}\left(\kappa^{\mathrm{ems}}\cdot{\mathbf{W}}_{t}^{T}{\mathbf{h}}_{t,n}\right),\end{split}$ (9)
which is equivalent to a softmax with temperature-scaled logits, with the temperature set to $1/\kappa^{\mathrm{ems}}$. Temperature scaling only affects the probabilities, not the modal class prediction. If using class-specific $\kappa^{\mathrm{ems}}$ values and assuming imbalanced classes, then these terms show up as class-specific bias terms: $p({\textnormal{c}}_{t,n,k}=1|{\mathbf{h}}_{t,n},{\mathbf{w}}_{t,1},\ldots,{\mathbf{w}}_{t,K},\kappa_{1}^{\mathrm{ems}},\ldots,\kappa_{K}^{\mathrm{ems}})\propto\exp\left\{\kappa_{k}^{\mathrm{ems}}\cdot{\mathbf{w}}_{t,k}^{T}{\mathbf{h}}_{t,n}+\log C_{D}(\kappa_{k}^{\mathrm{ems}})+\log\pi_{t,k}\right\}$ where $C_{D}(\kappa_{k}^{\mathrm{ems}})$ is the vMF’s normalization constant and $\pi_{t,k}$ is the mixing weight.
@@ 4 Related Work
State-Space Models for Deep Learning
@@
Probabilistic state-space models, and the Kalman filter (Kalman, 1960) in particular, have found diverse use in deep learning as a principled way of updating a latent state with new information. In sequence modelling, filter-based architectures are used to learn the latent state of a trajectory in both discrete (Krishnan et al., 2015; Karl et al., 2017; Fraccaro et al., 2017; Becker et al., 2019) and continuous time (Schirmer et al., 2022; Ansari et al., 2023; Zhu et al., 2023). Inspired by traditional state-space models, recent work on structured state-space models (Gu et al., 2022; Smith et al., 2023; Gu & Dao, 2023) has pushed the state-of-the-art in sequence modeling. However, in this setting, the SSM models the dynamics for an individual sequence while we are interested in modeling the dynamics of the data stream as a whole. The latter is also the subject of supervised online learning, and Chang et al. (2023) and Titsias et al. (2023) employ Kalman filters in the supervised, non-stationary setting. Similarly to our method, Titsias et al. (2023) also infer with an SSM the evolution of the linear classification head. However, they rely on labels to update the latent state of the weights whereas our weight adaptation is entirely unsupervised.
Test-Time Adaptation
TTA improves model performance on unseen test data with distribution shifts by adapting to the test data during inference. Early TTA approaches focused on recalculating batch normalization (BN) statistics from test data (Nado et al., 2020; Schneider et al., 2020). This has often been combined with minimizing entropy (Liang et al., 2020; Wang et al., 2021; Zhang et al., 2022; Yu et al., 2024; Gao et al., 2024). Alternative objectives leverage contrastive learning (Chen et al., 2022), invariance regularization (Nguyen et al., 2023), and Hebbian learning (Tang et al., 2023). While many of these methods can be applied online, they don’t fully address the challenges of non-stationary test streams. CTTA aims to handle changing distributions without forgetting source knowledge, using strategies such as episodic resets (Press et al., 2024), student-teacher models (Wang et al., 2022; Döbler et al., 2023; Brahma & Rai, 2023), masking (Liu et al., 2024), and regularization (Niu et al., 2022; Song et al., 2023). However, temporally correlated data remains challenging as they can cause online class imbalance and disrupt BN statistics, leading to model collapse. Solutions include adapted BN strategies (Zhao et al., 2023a; Lim et al., 2023), reservoir sampling (Gong et al., 2022; Yuan et al., 2023), filtering unreliable samples (Niu et al., 2023), and tracking label distributions (Zhou et al., 2023). Methods that bypass BN and instead adapt the classification head avoid collapse under temporal correlation. These methods adjust logits (Boudiaf et al., 2022) and predict based on KNNs (Jang et al., 2023). Most similar to our method, T3A (Iwasawa & Matsuo, 2021) recomputes prototypes from representations. However, T3A relies upon heuristics while STAD explicitly models dynamics using a SSM.

### v3

@@ 3.2 A Probabilistic Model of Shift Dynamics
where $\phi$ comprises the parameters of the transition and emission densities as well as the mixing coefficients, $\phi=\{\psi^{\mathrm{trans}},\psi^{\mathrm{ems}},\bm{\pi}_{1:T}\}$, and we have abbreviated the posterior distribution by $p({\mathbf{W}},{\mathbf{C}}\mid{\mathbf{H}}):=p({\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T}|{\mathbf{H}}_{1:T})$. After one optimization step, we collect the $K$ class prototypes into a matrix ${\mathbf{W}}_{t}$. Using the same hidden representations used to fit ${\mathbf{W}}_{t}$, we generate the predictions via the model’s softmax parameterization,
$y_{t,n}\ \sim\ \text{{Cat}}\left({\textnormal{y}}_{t,n};\ \text{{softmax}}({\mathbf{W}}_{t}^{\top}{\mathbf{h}}_{t,n})\right),$ (5)
where $y_{t,n}$ denotes a prediction sampled for the representation vector ${\mathbf{h}}_{t,n}$. Note that adaptation can be performed online by optimizing Eqn.˜4 incrementally, considering only data up to point $t$. To omit computing the complete-data log likelihood for an increasing sequence as time goes on, we employ a sliding window approach. Algorithm˜1 outlines the overall, procedure of our method. The specific EM updates depend on the chosen parametric form of the SSM. We consider two instances: a Gaussian model and a hyperspherical model based on the von Mises–Fisher distribution. The corresponding EM steps for these two cases are detailed in Algorithm˜2 and Algorithm˜3, respectively.
Algorithm 1 STAD
1:Input: Source model $f_{\theta}$, test batches ${\mathbf{X}}_{1:T}$, sliding window size $s$
2:Initialize: mixing coefficients $\bm{\pi}_{t}$, weights ${\mathbf{W}}_{t}$, transition and emission parameters $\psi^{\text{trans}},\psi^{\text{ems}}$
3:for $t\in\mathcal{T}$ do
4:  Define sliding window: $S_{t}=\{\tau\mid\max(1,t-s)\leq\tau\leq t\}$
5:  Compute representations: ${\mathbf{H}}_{t}=f_{\theta}^{L-1}({\mathbf{X}}_{t})$
6:  Fit probabilistic SSM via EM: ${\mathbf{W}}_{t},{\mathbf{C}}_{t},\bm{\pi}_{t},\psi^{\text{trans}},\psi^{\text{ems}}=\text{EM}\big(\{{\mathbf{H}}_{\tau},{\mathbf{W}}_{\tau},{\mathbf{C}}_{\tau},\bm{\pi}_{\tau}\}_{\tau\in S_{t}},\psi^{\text{trans}},\psi^{\text{ems}}\big)$
7:  Predict: $y_{t,n}\sim\ \text{{Cat}}\big({\textnormal{y}}_{t,n};\ \text{{softmax}}({\mathbf{W}}_{t}^{\top}{\mathbf{h}}_{t,n})\big)$
8:end for
@@
The simplest parametric choice for the transition and emission models is Gaussian. The resulting probabilistic SSM can be interpreted as a mixture of $K$ Kalman filters (KFs) (Kalman, 1960). Owing to the linear–Gaussian structure, the posterior expectation $\mathbb{E}_{p({\mathbf{W}},{\mathbf{C}}|{\mathbf{H}})}[\cdot]$ in Eqn.˜4 can be computed in closed form using the standard KF prediction, update, and smoothing equations (Calabrese & Paninski, 2011; Bishop & Nasrabadi, 2006). The complete model specification is provided in Sec.˜C.1, and the corresponding EM updates are summarized in Algorithm˜2. However, these closed-form computations come at a cost: they require matrix inversions of size $D\times D$ and the storage of $K\times D^{2}$ parameters. This makes the Gaussian formulation expensive for high-dimensional features and impractical in low-resource settings. Next, we present a model for spherical features that circumvents these limitations.
@@ 3.3 Von Mises-Fisher Model for Hyperspherical Features
Choosing Gaussian densities for the transition and emission models, as discussed above, assumes the representation space follows an Euclidean geometry. However, prior work has shown that assuming the hidden representations lie on the unit hypersphere results in a better inductive bias for OOD generalization (Mettes et al., 2019; Bai et al., 2024). This is due to the norms of the representations being biased by in-domain information such as class balance, making angular distances a more reliable signal of class membership in the presence of distribution shift (Mettes et al., 2019; Bai et al., 2024). We too employ the hyperspherical assumption by normalizing the hidden representations such that ${\mathbf{h}}||_{2}=1$ and model them with the von Mises-Fisher (vMF) distribution (Mardia & Jupp, 2009),
${\text{{vMF}}}({\mathbf{h}};\bm{\mu}_{k},\kappa)=C_{D}(\kappa)\exp\left\{\kappa\cdot\bm{\mu}_{k}^{\top}{\mathbf{h}}\right\}$ (6)
where $\bm{\mu}_{k}\in{\mathbb{R}}^{D}$ with $\bm{\mu}_{k}||_{2}=1$ denotes the mean direction of class $k$, $\kappa\in{\mathbb{R}}^{+}$ the concentration parameter, and $C_{D}(\kappa)$ the normalization constant. High values of $\kappa$ imply larger concentration around $\bm{\mu}_{k}$. The vMF distribution is proportional to a Gaussian distribution with isotropic variance and unit norm. While previous work (Mettes et al., 2019; Ming et al., 2023; Bai et al., 2024) has explored training objectives to encourage representations to be vMF-distributed, we apply Eqn.˜6 to model the evolving representations.
@@
Returning to the SSM given above (Eqn.˜3), we specify both transition (Eqn.˜1) and emission models (Eqn.˜2) as vMF distributions, resulting in a hyperspherical transition model, $p({\mathbf{W}}_{t}|{\mathbf{W}}_{t-1})=\prod_{k=1}^{K}{\text{{vMF}}}({\mathbf{w}}_{t,k}|{\mathbf{w}}_{t-1,k},\kappa^{\mathrm{trans}})$, and hyperspherical emission model, $p({\mathbf{H}}_{t}|{\mathbf{W}}_{t})=\prod_{n=1}^{N_{t}}\sum_{k=1}^{K}\pi_{t,k}{\text{{vMF}}}({\mathbf{h}}_{t,n}|{\mathbf{w}}_{t,k},\kappa^{\mathrm{ems}})$. The parameter size of the vMF formulation scales linearly with the feature dimension, $\mathcal{O}(DK)$, compared to the Gaussian case’s $\mathcal{O}(D^{2}K)$. Notably, the noise parameters, $\kappa^{\mathrm{trans}},\kappa^{\mathrm{ems}}$ simplify to scalar values which reduces memory substantially. Fig.˜2 illustrates this STAD-vMF variant.
@@
In contrast to the linear Gaussian case, the vMF distribution is not closed under marginalization. As a result, the posterior distribution required for computing the expectation in Eqn.˜4 (E-step of the EM algorithm) cannot be expressed in closed form. To address this, we adopt a variational EM approach, approximating the posterior $p({\mathbf{W}},{\mathbf{C}}\mid{\mathbf{H}})$ with a mean-field variational distribution $q({\mathbf{W}},{\mathbf{C}})$ following Gopal & Yang (2014):
$q({\mathbf{w}}_{t,k})={\text{{vMF}}}(\cdot;\bm{\rho}_{t,k},\gamma_{t,k}),\quad q({\mathbf{c}}_{t,n})=\text{{Cat}}(\cdot;\bm{\lambda}_{t,n}).$ (7)
The variational distribution $q({\mathbf{W}},{\mathbf{C}})$ factorizes over $t,n,k$ and the objective from Eqn.˜4 becomes $\operatorname*{arg\,max}_{\phi}\mathbb{E}_{q({\mathbf{W}},{\mathbf{C}})}\big[\log p({\mathbf{H}}_{1:T},{\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T})\big]$. The optimal variational parameters are given by
[…]
$\text{with}\quad\beta_{t,k}$ $=\kappa^{\mathrm{trans}}\,\mathbb{I}_{\{t>1\}}\,\mathbb{E}_{q}[{\mathbf{w}}_{t-1,k}]+\kappa^{\mathrm{ems}}\sum_{n=1}^{N_{t}}\mathbb{E}_{q}[c_{t,n,k}]{\mathbf{h}}_{t,n}+\kappa^{\mathrm{trans}}\,\mathbb{I}_{\{t<T\}}\,\mathbb{E}_{q}[{\mathbf{w}}_{t+1,k}],$ (9)
where $\mathbb{I}_{\{\cdot\}}$ denotes the indicator function, ensuring that terms are omitted at the temporal boundaries and $\mathbb{E}_{q}$ denotes expectations with respect to $q({\mathbf{W}},{\mathbf{C}})$. The expectations of the variational distributions (Eqn.˜7) are $\mathbb{E}_{q}[{\mathbf{c}}_{t,n}]=(\lambda_{t,n,1},\dots,\lambda_{t,n,K})$ and $\mathbb{E}_{q}[{\mathbf{w}}_{t,k}]=A_{D}(\gamma_{t,k})\bm{\rho}_{t,k}$, which completes the E-step. We defer the M-step to Sec.˜C.2. Algorithm˜3 summarizes the EM algorithm for STAD-vMF. As in the Gaussian case, we form the updated last-layer weight matrix by stacking the posterior means, ${\mathbf{W}}_{t}=(\bm{\rho}_{t,1},\dots,\bm{\rho}_{t,K})$. Notably, posterior inference for the vMF model is much more scalable than the Gaussian case. It operates with linear complexity in $D$, rather than cubic, reducing runtime significantly.
@@
$\begin{split}p&({\textnormal{c}}_{t,n,k}=1|{\mathbf{h}}_{t,n},{\mathbf{w}}_{t,1},\ldots,{\mathbf{w}}_{t,K},\kappa^{\mathrm{ems}})\ =\ \frac{{\text{{vMF}}}({\mathbf{h}}_{t,n};{\mathbf{w}}_{t,k},\kappa^{\mathrm{ems}})}{\sum_{j=1}^{K}{\text{{vMF}}}({\mathbf{h}}_{t,n};{\mathbf{w}}_{t,j},\kappa^{\mathrm{ems}})}\\ &=\frac{C_{D}(\kappa^{\mathrm{ems}})\exp\left\{\kappa^{\mathrm{ems}}\cdot{\mathbf{w}}_{t,k}^{\top}{\mathbf{h}}_{t,n}\right\}}{\sum_{j=1}^{K}C_{D}(\kappa^{\mathrm{ems}})\exp\left\{\kappa^{\mathrm{ems}}\cdot{\mathbf{w}}_{t,j}^{\top}{\mathbf{h}}_{t,n}\right\}}\ =\ \text{{softmax}}\left(\kappa^{\mathrm{ems}}\cdot{\mathbf{W}}_{t}^{\top}{\mathbf{h}}_{t,n}\right),\end{split}$ (10)
which is equivalent to a softmax with temperature-scaled logits, with the temperature set to $1/\kappa^{\mathrm{ems}}$. Temperature scaling only affects the probabilities, not the modal class prediction. If using class-specific $\kappa^{\mathrm{ems}}$ values and assuming imbalanced classes, then these terms show up as class-specific bias terms, $p({\textnormal{c}}_{t,n,k}=1|{\mathbf{h}}_{t,n},{\mathbf{w}}_{t,1},\ldots,{\mathbf{w}}_{t,K},\kappa_{1}^{\mathrm{ems}},\ldots,\kappa_{K}^{\mathrm{ems}})\propto\exp\left\{\kappa_{k}^{\mathrm{ems}}\cdot{\mathbf{w}}_{t,k}^{\top}{\mathbf{h}}_{t,n}+\log C_{D}(\kappa_{k}^{\mathrm{ems}})+\log\pi_{t,k}\right\}$.
@@ 4 Related Work
We overview SSMs and TTA next. App.˜B provides more detailed discussions on TTA and adjacent fields.
State-Space Models (SSMs) in Deep Learning
@@
Probabilistic SSMs, such as the Kalman filter (Kalman, 1960), provide a principled framework for updating latent states with new information and have been widely applied in deep learning. In sequence modeling, SSMs are used to learn latent trajectories in both discrete (Krishnan et al., 2015; Karl et al., 2017; Fraccaro et al., 2017; Becker et al., 2019) and continuous time (Schirmer et al., 2022; Ansari et al., 2023; Zhu et al., 2023). Recent advancements in structured SSMs (Gu et al., 2022; Smith et al., 2023; Gu & Dao, 2023) have pushed the state-of-the-art in sequence modeling. However, these models focus on individual sequence dynamics, whereas we are interested in modeling the dynamics of an entire data stream. Our objective aligns more closely with online learning (Duran-Martin et al., 2025). Notably, Chang et al. (2023) and Titsias et al. (2024) extend Kalman filters to handle non-stationary, supervised settings. Like our method, Titsias et al. (2024) infers the evolution of the classification head with a SSM. However, they require labels whereas our method is fully unsupervised.
Test-Time Adaptation (TTA)
TTA aims to make pre-trained models robust to distribution shifts by adapting directly to the test data during inference. It is training-agnostic, contributing to its growing popularity (Xiao & Snoek, 2024). Early TTA methods recalculate batch normalization (BN) statistics from test data (Nado et al., 2020; Schneider et al., 2020). Updating model parameters via gradient descent is most commonly done via entropy minimization (Grandvalet & Bengio, 2004; Wang et al., 2021; Zhang et al., 2022; Yu et al., 2024; Gao et al., 2024). Alternative approaches include contrastive learning (Chen et al., 2022), invariance regularization (Nguyen et al., 2023), Hebbian learning (Tang et al., 2023), and prompt tuning (Niu et al., 2024). Recent work has focused on making TTA reliably deployable by stress-testing various real-world settings (see App.˜B for detailed discussions). A key challenge is continual adaptation to a non-stationary target domain, studied in CTTA (Wang et al., 2022b). Solutions include episodic resets (Press et al., 2024), student-teacher models (Döbler et al., 2023; Brahma & Rai, 2023), masking (Liu et al., 2024) and regularization (Niu et al., 2022; Song et al., 2023). Approaches relying on test data statistics further struggle with class imbalance or small test batches, which has led to adaptations in BN strategies (Zhao et al., 2023a; Lim et al., 2023), reservoir sampling (Gong et al., 2022; Yuan et al., 2023), sample filtering (Niu et al., 2023), and label distribution tracking (Zhou et al., 2023). Methods that adapt the classification head, rather than relying on BN, effectively prevent such collapse (Boudiaf et al., 2022; Jang et al., 2023). The most similar approach to ours, T3A (Iwasawa & Matsuo, 2021), recomputes prototypes from representations but relies on heuristics, whereas STAD explicitly models dynamics with a SSM. Lee & Chang (2024b; a); Lee (2025) also use a SSM for online TTA, but to filter noisy model updates, while our SSM models the distribution shift itself. Concurrent to our work, Dai & Yang (2025) proposes Gaussian mixture models updated via EM to adapt the prototypes of vision language models (VLMs). Aside from focusing on VLMs, their model is also static and does not consider transition dynamics.